# Decentralized Nonconvex Optimization under Heavy-Tailed Noise: Normalization and Optimal Convergence

### v1

@@ Abstract
Heavy-tailed noise in nonconvex stochastic optimization has garnered increasing research interest, as empirical studies, including those on training attention models, suggest it is a more realistic gradient noise condition. This paper studies first-order nonconvex stochastic optimization under heavy-tailed gradient noise in a decentralized setup, where each node can only communicate with its direct neighbors in a predefined graph. Specifically, we consider a class of heavy-tailed gradient noise that is zero-mean and has only $p$-th moment for $p\in(1,2]$. We propose GT-NSGDm, Gradient Tracking based Normalized Stochastic Gradient Descent with momentum, that utilizes normalization, in conjunction with gradient tracking and momentum, to cope with heavy-tailed noise on distributed nodes. We show that, when the communication graph admits primitive and doubly stochastic weights, GT-NSGDm guarantees, for the first time in the literature, that the expected gradient norm converges at an optimal non-asymptotic rate $O\big{(}1/T^{(p-1)/(3p-2)}\big{)}$, which matches the lower bound in the centralized setup. When tail index $p$ is unknown, GT-NSGDm attains a non-asymptotic rate $O\big{(}1/T^{(p-1)/(2p)}\big{)}$ that is, for $p<2$, topology independent and has a speedup factor $n^{1-1/p}$ in terms of the number of nodes $n$. Finally, experiments on nonconvex linear regression with tokenized synthetic data and decentralized training of language models on a real-world corpus demonstrate that GT-NSGDm is more robust and efficient than baselines.
@@ 1 Introduction
In this paper, we address the problem of nonconvex stochastic optimization under heavy-tailed gradient noise in the decentralized setup. Consider a graph with $n$ nodes connected by a predefined topology $\mathcal{G}:=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{1,\ldots,n\}$ is the set of node indices, and $\mathcal{E}$ is the collection of directed pairs $(i,r)$, $i,r\in\mathcal{V}$ such that node $i$ can send information to the neighboring node $r$. Each node $i\in\mathcal{V}$ holds a local nonconvex differentiable cost function $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and can access its stochastic gradient, subject to zero mean noise with a bounded $p$-th moment for some $p\in(1,2]$. Cooperatively, these nodes aim to solve $\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x}):=(1/n)\sum_{i=1}^{n}f_{i}(\bm{x})$, through local computation and peer-to-peer communication only with their immediate neighbors,
@@
Decentralized optimization under heavy-tailed gradient noise remains underexplored. To the best of our knowledge, only recent studies [29, 30] have attempted to address this problem under restrictive assumptions. Specifically, [29] considers zero-mean gradient noise with bounded $p$-th central moment ($p\in(1,2]$) similar to our setting but assumes a compact domain, effectively implying bounded gradients. Their proposed decentralized gradient descent method with $\ell_{2}$ gradient clipping achieves almost sure convergence for strongly convex local functions. However, the restrictive compact domain assumption in [29] limits its practical applicability, and the convergence rate is not explicitly provided. Another work, [30], also assumes strongly convex local objectives and develops a decentralized gradient method with smoothed clipping and error feedback under gradient noise that is zero-mean, symmetric, and has bounded first absolute moment, showing an in-expectation convergence rate of $1/t^{\delta}$ for some $\delta\in(0,2/5)$. Although the noise assumption in [30] is weaker than ours (as it requires only a first-moment bound), the additional assumptions of noise symmetry and the dependence of the rate exponent $\delta$ on both the problem dimension and condition number restrict its general applicability. Moreover, both works [29, 30] assume strong convexity, whereas many practical optimization problems involving heavy-tailed noise, particularly in modern machine learning, are inherently nonconvex. Further, the convergence rates in [29, 30] are either unclear or sub-optimal, even compared to the optimal iteration complexity bound $O\big{(}1/T^{(p-1)/(3p-2)}\big{)}$ for general nonconvex functions. In this work, we relax these restrictive assumptions and address the following question:
@@ 1.1 Contributions
We provide a positive answer to the above question through the following key contributions.
  • We develop a decentralized method, called GT-NSGDm, using normalization, coupled with momentum variance reduction, to combat heavy-tailed noise, and using gradient tracking to handle cross-node heterogeneity. To further shed light on the design of GT-NSGDm, we provide a negative result for a vanilla variant of normalized decentralized SGD that employs no gradient tracking nor momentum.
  • For general nonconvex and smooth local functions $f_{i}$’s that are bounded from below, we show that GT-NSGDm converges in expectation at a rate $O\big{(}1/T^{(p-1)/(3p-2)}\big{)}$, which matches the lower bound in centralized setting and is order-optimal. Our convergence rate significantly improves upon related works [29, 30], which assume strong convexity and either do not provide a convergence rate or lack an explicit rate exponent.
@@ 1.2 Related Work
Adaptive methods. Under heavy-tailed noise, vanilla SGD based methods are shown to suffer from slower convergence or even model collapses in centralized settings [17] as well as distributed settings with a central server [23, 35], and adaptive methods such as clipping and normalization are introduced to stabilize training dynamics. In centralized settings, the work [17] provides lower bounds for both nonconvex and strongly convex smooth functions, showing that SGD with gradient clipping achieves in-expectation upper bounds matching lower bounds. In [24, 36, 37, 38], the authors show that when equipped with gradient clipping, SGD, accelerated methods, AdaGrad [39], and Adam [40] can achieve (near-)optimal high-probability convergence under various function assumptions. Besides, the work [25] shows that signSGD is also robust to heavy-tailed noise through the lens of stochastic differential equations. Further, SGD with gradient normalization, which advantageously requires less hyper-parameter tuning than clipping, is shown to achieve optimal in-expectation convergence [26, 27, 41]. Our method incorporates the same normalization and variance reduction approach as [27]. Notably, in another line of works [42, 28, 43], the authors conduct a unified convergence analyses for generic nonlinear methods including clipping, sign, and normalization under symmetric noise with positive probability mass around zero without assuming any noise moment bound or only assuming first absolute noise moment bound. In distributed settings with a server, the work [44] proposes an algorithm that incorporates an error feedback mechanism, wherein clipping is applied to the discrepancy between a local gradient estimator and a stochastic gradient, and establishes optimal high-probability bounds. Moreover, the work [45] shows that distributed signSGD converges to an asymptotic neighborhood depending on the ‘fatness’ of noise tail. When multiple local updates are permitted between communication rounds, the authors of [23] show that clipping per local step achieves order-optimal in-expectation convergence, albeit under a restrictive bounded gradient assumption. More recently, the work [35] introduces the TailOPT framework, which adaptively leverages gradient geometry by applying clipping operators during local updates on distributed nodes and utilizing adaptive optimizers for global updates at the server, achieving in-expectation sublinear convergence rates that are independent of the moment parameter $p$.
@@
Nonlinearities in decentralized optimization. Extending existing methods that are robust to heavy-tailed noise, whether developed for centralized settings or distributed settings with a server, to decentralized environments is highly nontrivial, primarily due to the nonlinearities introduced to peer-to-peer communication. This difficulty is reflected in that existing decentralized methods incorporating nonlinear adaptive techniques for other purposes often impose restrictive conditions [46, 47]. For example, to achieve differential privacy through gradient clipping, the work [47] establishes convergence in decentralized setups under the assumption of either bounded gradient or a stringent similarity condition, namely $\|\nabla f_{i}(\bm{x})-\nabla f(\bm{x})\|\leq(1/12)\|\nabla f(\bm{x})\|$ for all $i\in[n]$ and all $\bm{x}$. Similarly, to attain adversarial robustness against gradient attacks, the authors of [46] employ gradient clipping with momentum, assuming that all local functions are convex, share a common minimizer, and that $\sum_{i=1}^{n}f_{i}$ is strongly convex. In this work, we significantly relax these conditions and demonstrate the effective use of nonlinearity (specifically, normalization) in decentralized optimization, thereby motivating broader applications of nonlinear techniques in this setting.
1.3 Paper Outline and Notation
We introduce the problem setup and develop our proposed algorithm GT-NSGDm in Sections 2 and 3, respectively. Then, we present our convergence results in Section 4 and discuss their implications. Section 5 presents our numerical studies, while Section 6 concludes the paper.
@@ 1.3 Paper Outline and Notation
We denote by $\mathbb{N}_{+}$, $\mathbb{R}$, $\mathbb{R}_{+}$ and $\mathbb{R}^{d}$, respectively, the set of positive natural numbers, real numbers, nonegative real numbers, and the $d$-dimensional Euclidean space. We use lowercase normal letters for scalars, lowercase boldface letters for vectors, and uppercase boldface letters for matrices. Further, we denote by $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ the all-ones and all-zeros vectors of size $k$, respectively, and by $\bm{I}_{k}$ the $k\times k$ identity matrix. We let $\|\bm{x}\|$ denote the $\ell_{2}$ norm of $\bm{x}$, and $\|\bm{A}\|_{2}$ denote the operator norm of $\bm{A}$. For functions $p(t)$ and $q(t)$ in $t$, we write $p(t)=O(q(t))$ if $\limsup_{t\to\infty}p(t)/q(t)<\infty$. Finally, we use $\mathbb{E}$ to denote expectation over over random quantities.
@@ 2 Problem Formulation
We consider a graph with $n$ nodes, where each node holds a local and private function $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and the nodes collectively minimize the unconstrained global objective $f(\bm{x}):=(1/n)\sum_{i=1}^{n}f_{i}(\bm{x})$ through peer-to-peer communication. We now present the assumptions on the problem model.
@@
The above assumptions are standard in decentralized optimization, encompassing both empirical risk minimization and online optimization settings. Examples of local functions satisfying these assumptions include regularized linear regression, logistic regression, and neural networks [2, 48].
@@
We then assume this stochastic first-order oracle have the following properties.
@@
For any $\mathcal{F}_{t}$-measurable random vectors $\bm{x}\in\mathbb{R}^{d}$, we have the following: $\forall i\in[n],\forall t\geq 0$,
  • $\mathbb{E}[\bm{g}_{i}(\bm{x},\bm{\xi}_{i}^{t})\mid\mathcal{F}_{t}]=\nabla f_{i}(\bm{x})$;
  • There exist $p\in(1,2]$, some constant $\sigma\geq 0$ such that $\mathbb{E}\big{[}\|\bm{g}_{i}(\bm{x},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x})\|^{p}\mid\mathcal{F}_{t}\big{]}\leq\sigma^{p}$;
  • The family $\{\bm{\xi}_{i}^{t}:\forall t\geq 0,i\in[n]\}$ of random samples is independent.
@@
Assumption 3 covers a broad class of heavy-tailed distributions, including Lévy’s $\alpha$-stable distributions, Student’s $t$-distributions, and Pareto distributions. Note that we do not assume noise symmetry as in [30], and when $p=2$, Assumption 3 reduces to the standard bounded variance condition commonly assumed in the literature [49].
@@
The nonnegative weight matrix $\bm{W}$, whose $(i,r)$-th component of $\bm{W}$, denoted as $w_{ir}$, is positive if and only if $(i,r)\in E$ or $i=r$, has the following properties:
  • $\bm{W}$ is primitive;
  • $\bm{W}$ is doubly stochastic, i.e., $\mathbf{1}_{n}\bm{W}=\mathbf{1}_{n}$ and $\bm{W}\mathbf{1}_{n}=\mathbf{1}_{n}$.
Assumption 4 is standard in the decentralized optimization literature [51], and it guarantees that there exits some nonnegative $\lambda$, i.e., spectral gap, such that
@@ 3 Algorithm Development: GT-NSGDm
We now describe the proposed Algorithm GT-NSGDm and discuss the intuition of its construction. We use $\bm{x}_{i}^{t}$ to denote the estimate of a stationary point for the global cost function $f$ at node $i$ and $t$-th iteration, and recall that $\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})$ denotes the corresponding stochastic gradient returned from local first-order oracle. Motivated by the error-feedback approach in [30], which serves as a momentum-type of variance reduction after applying a nonlinear operator to handle heavy-tailed noise, we also employ local momentum variance reduction
@@
where $\beta\in[0,1)$ serves as the momentum coefficient. Then, we use gradient tracking [12] to handle heterogeneous local functions $\{f_{i}\}_{i=1}^{n}$. Specifically, we use an estimator $\bm{y}_{i}^{t}$ to track global gradient
@@
It is known that gradient tracking [12] helps eliminate the dependence on heterogeneity among local functions $\{f_{i}\}_{i=1}^{n}$, such as the requirement of bounded gradient similarity. Furthermore, similar to the approach in [27], which uses normalization to address heavy-tailed noise in centralized settings, we avoid applying normalization in the recursive updates of the local gradient estimator $\bm{v}_{i}^{t}$ in (1) and the global gradient tracker $\bm{y}_{i}^{t}$. Instead, normalization is applied only during the update of $\bm{x}_{i}^{t}$, with step size $\alpha$, and nonnegative mixing weights $\{w_{ir}\}$ that $w_{ir}>0$ only when $(i,r)\in\mathcal{E}$ or $i=r$,
@@
We combine the local updates (1)(2)(3) on node $i\in\mathcal{V}$ and call it GT-NSGDm, Gradient Tracking based Normalized Stochastic Gradient Descent with momentum. When taking $\beta=0$, this simplifies to momentum-free gradient tracking with normalization in step 3. However, our analysis shows that GT-NSGDm performs optimal for some $\beta\in(0,1)$, making GT-NSGDm a non-trivial and optimal algorithmic design for the considered problem class. We provide a tabular description for GT-NSGDm in Algorithm 1, where all $\{\bm{x}_{i}^{0}\}$ are initialized from the same point $\bar{\bm{x}}^{0}$ for simplicity.
@@
Then, the global average $\bar{\bm{x}}^{t}$ would update in the negative direction of the sum of normalized local gradients: $\bar{\bm{x}}^{t+1}=\bar{\bm{x}}^{t}-\frac{\alpha}{n}\sum_{r=1}^{n}\frac{\nabla f_{r}(\bm{x}_{r}^{t})}{\|\nabla f_{r}(\bm{x}_{r}^{t})\|}.$ Let for some $t,\forall r\in\mathcal{V},\bm{x}_{r}^{t}=\bm{x}_{*}=\mathop{\mathrm{arg\,min}}\sum_{r=1}^{n}f_{i}(\bm{x})$, i.e., all nodes hold the optimal global solution that $\sum_{r=1}^{n}\nabla f_{r}(\bm{x}_{*})=\mathbf{0}$. Since $\|\nabla f_{r}(\bm{x}_{*})\|$ can be different quantities for $r=1,\ldots,n$, due to function heterogeneity, then $\bar{\bm{x}}^{t+1}$ will move away from $\bm{x}_{*}$. Therefore, vanilla gradient normalization adds some intrinsic errors from heterogeneous local normalizations. By incorporating gradient tracking, we expect that $\bm{y}_{r}^{t}$ would converge to its global average $\bar{\bm{y}}^{t}$, and $\bar{\bm{y}}^{t}$ would converge to $(1/n)\sum_{i=r}^{n}\nabla f_{i}(\bm{x}_{r}^{t})$. In this way, $\bm{x}_{r}^{t}$ would move along the direction of the normalized sum of local gradients, and thus emulating the centralized setting.
@@
We break the limitations of vanilla gradient normalization in Claim 1 by incorporating gradient tracking and momentum-based variance reduction. This enables the successful use of normalization to suppress heavy-tailed noise while maintaining optimal convergence despite the added nonlinearity.

### v2

@@ Abstract
Heavy-tailed noise in nonconvex stochastic optimization has garnered increasing research interest, as empirical studies, including those on training attention models, suggest it is a more realistic gradient noise condition. This paper studies first-order nonconvex stochastic optimization under heavy-tailed gradient noise in a decentralized setup, where each node can only communicate with its direct neighbors in a predefined graph. Specifically, we consider a class of heavy-tailed gradient noise that is zero-mean and has only $p$-th moment for $p\in(1,2]$. We propose GT-NSGDm, Gradient Tracking based Normalized Stochastic Gradient Descent with momentum, that utilizes normalization, in conjunction with gradient tracking and momentum, to cope with heavy-tailed noise on distributed nodes. We show that, when the communication graph admits primitive and doubly stochastic weights, GT-NSGDm guarantees, for the first time in the literature, that the expected gradient norm converges at an optimal non-asymptotic rate $O\big(1/T^{(p-1)/(3p-2)}\big)$, which matches the lower bound in the centralized setup. When the tail index $p$ is unknown, GT-NSGDm attains a non-asymptotic rate $O\big(1/T^{(p-1)/(2p)}\big)$ that is, for $p<2$, topology independent and has a speedup factor $n^{1-1/p}$ in terms of the number of nodes $n$. Finally, experiments on nonconvex linear regression with tokenized synthetic data and decentralized training of language models on a real-world corpus demonstrate that GT-NSGDm is more robust and efficient than baselines.
@@ 1 Introduction
In this paper, we address the problem of nonconvex stochastic optimization under heavy-tailed gradient noise in the decentralized setup. Consider a graph with $n$ nodes connected by a predefined topology $\mathcal{G}:=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{1,\ldots,n\}$ is the set of node indices, and $\mathcal{E}$ is the collection of directed pairs $(i,r)$, $i,r\in\mathcal{V}$ such that node $i$ can send information to the neighboring node $r$. Each node $i\in\mathcal{V}$ holds a local nonconvex differentiable cost function $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and can access its stochastic gradient, subject to zero mean noise with a bounded $p$-th moment for some $p\in(1,2]$. Cooperatively, these nodes aim to solve $\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x}):=(1/n)\sum_{i=1}^{n}f_{i}(\bm{x})$, through local computation and peer-to-peer communication only with their immediate neighbors.
@@
Decentralized optimization under heavy-tailed gradient noise remains underexplored. To the best of our knowledge, only recent studies [29, 30] have attempted to address this problem under restrictive assumptions. Specifically, [29] considers zero-mean gradient noise with bounded $p$-th central moment ($p\in(1,2]$) similar to our setting but assumes a compact domain or bounded gradients. Their proposed decentralized gradient descent method with $\ell_{2}$ gradient clipping achieves almost sure convergence for strongly convex local functions. However, the restrictive compact domain or gradients assumption in [29] limits its practical applicability, and the convergence rate is not explicitly provided. Another work, [30], also assumes strongly convex local objectives and develops a decentralized gradient method with smoothed clipping and error feedback under gradient noise that is zero-mean, symmetric, and has bounded first absolute moment, showing an in-expectation convergence rate of $1/t^{\delta}$ for some $\delta\in(0,2/5)$. Although the noise assumption in [30] is weaker than ours (as it requires only a first-moment bound), the additional assumptions of noise symmetry and the dependence of the rate exponent $\delta$ on both the problem dimension and condition number restrict its general applicability. Moreover, both works [29, 30] assume strong convexity, whereas many practical optimization problems involving heavy-tailed noise, particularly in modern machine learning, are inherently nonconvex. Further, the convergence rates in [29, 30] are either unclear or sub-optimal, even compared to the optimal iteration complexity bound $O\big(1/T^{(p-1)/(3p-2)}\big)$ for general nonconvex functions. In this work, we relax these restrictive assumptions and address the following question:
@@ 1.1 Contributions
We answer this question affirmatively through the following key contributions:
  • We develop a decentralized method, called GT-NSGDm, using normalization, coupled with momentum variance reduction, to combat heavy-tailed noise, and using gradient tracking to handle cross-node heterogeneity. To further shed light on the design of GT-NSGDm, we provide a negative result for a vanilla variant of normalized decentralized SGD that employs neither gradient tracking nor momentum.
  • For general nonconvex and smooth local functions $f_{i}$’s that are bounded from below, we show that GT-NSGDm converges in expectation at a rate $O\big(1/T^{(p-1)/(3p-2)}\big)$, which matches the lower bound in the centralized setting and is order-optimal. Our convergence rate significantly improves upon related works [29, 30], which assume strong convexity and lack an explicit rate exponent.
@@ 1.2 Related Work
Adaptive methods. Under heavy-tailed noise, vanilla SGD based methods are shown to suffer from slower convergence or even model collapses in centralized settings [17] as well as distributed settings with a central server [23, 35], and adaptive methods such as clipping and normalization are introduced to stabilize training dynamics. In centralized settings, the work [17] provides lower bounds for both nonconvex and strongly convex smooth functions, showing that SGD with gradient clipping achieves in-expectation upper bounds matching lower bounds. In [24, 36, 37, 38], the authors show that when equipped with gradient clipping, SGD, accelerated methods, AdaGrad [39], and Adam [40] can achieve (near-)optimal high-probability convergence under various function assumptions. Besides, the work [25] shows that signSGD is also robust to heavy-tailed noise through the lens of stochastic differential equations. Further, SGD with gradient normalization, which advantageously requires less hyper-parameter tuning than clipping, is shown to achieve optimal in-expectation convergence [26, 27, 41]. Our method incorporates the same normalization and variance reduction approach as [27]. Notably, in another line of works [42, 28, 43], the authors conduct a unified convergence analysis for generic nonlinear methods including clipping, sign, and normalization under symmetric noise with positive probability mass around zero without assuming any noise moment bound or only assuming a first absolute noise moment bound. In distributed settings with a server, the work [44] proposes an algorithm that incorporates an error feedback mechanism, wherein clipping is applied to the discrepancy between a local gradient estimator and a stochastic gradient, and establishes optimal high-probability bounds. Moreover, the work [45] shows that distributed signSGD converges to an asymptotic neighborhood depending on the ‘fatness’ of the noise tail. When multiple local updates are permitted between communication rounds, the authors of [23] show that clipping per local step achieves order-optimal in-expectation convergence, albeit under a restrictive bounded gradient assumption. In addition, the work [46] rigorously shows that distributed training under heavy-tailed noise can collapse with unbounded gradients, and shows that client-side adaptive methods can stabilize training with bounded regrets in an online learning framework. More recently, the work [35] introduces the TailOPT framework, which adaptively leverages gradient geometry by applying clipping operators during local updates on distributed nodes and utilizing adaptive optimizers for global updates at the server, achieving in-expectation sublinear convergence rates that are independent of the moment parameter $p$.
@@
Nonlinearities in decentralized optimization. Extending existing methods that are robust to heavy-tailed noise, whether developed for centralized settings or distributed settings with a server, to decentralized environments is highly nontrivial, primarily due to the nonlinearities introduced to peer-to-peer communication. This difficulty is reflected in that existing decentralized methods incorporating nonlinear adaptive techniques for other purposes often impose restrictive conditions [47, 48]. For example, to achieve differential privacy through gradient clipping, the work [48] establishes convergence in decentralized setups under the assumption of either bounded gradient or a stringent similarity condition, namely $\|\nabla f_{i}(\bm{x})-\nabla f(\bm{x})\|\leq(1/12)\|\nabla f(\bm{x})\|$ for all $i\in[n]$ and all $\bm{x}$. Similarly, to attain adversarial robustness against gradient attacks, the authors of [47] employ gradient clipping with momentum, assuming that all local functions are convex, share a common minimizer, and that $\sum_{i=1}^{n}f_{i}$ is strongly convex. Furthermore, [49] integrates gradient tracking with normalization to address a max-margin problem. While gradient tracking is employed to ensure similarity among local gradient estimators, the method guarantees convergence only to the direction of the optimal solution under deterministic gradients. In this work, we significantly relax these conditions and demonstrate the effective use of nonlinearity (specifically, normalization) in decentralized optimization, thereby motivating broader applications of nonlinear techniques in this setting.
1.3 Notation
@@
We denote by $\mathbb{N}_{+}$, $\mathbb{R}$, $\mathbb{R}_{+}$ and $\mathbb{R}^{d}$, respectively, the set of positive natural numbers, real numbers, nonnegative real numbers, and the $d$-dimensional Euclidean space. We use lowercase normal letters for scalars, lowercase boldface letters for vectors, and uppercase boldface letters for matrices. Further, we denote by $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ the all-ones and all-zeros vectors of size $k$, respectively, and by $\bm{I}_{k}$ the $k\times k$ identity matrix. We let $\|\bm{x}\|$ denote the $\ell_{2}$ norm of $\bm{x}$, and $\|\bm{A}\|_{2}$ denote the operator norm of $\bm{A}$. For functions $p(t)$ and $q(t)$ in $t$, we write $p(t)=O(q(t))$ if $\limsup_{t\to\infty}p(t)/q(t)<\infty$. Finally, we use $\mathbb{E}$ to denote expectation over random quantities.
@@ 2 Problem Formulation
We consider a graph with $n$ nodes, where each node holds a local and private function $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and the nodes collectively minimize the unconstrained global objective $f(\bm{x}):=(1/n)\sum_{i=1}^{n}f_{i}(\bm{x})$ through peer-to-peer communication. We now present some standard assumptions on the problem.
@@
We then assume this stochastic first-order oracle has the following properties.
@@
For any $\mathcal{F}_{t}$-measurable random vectors $\bm{x}\in\mathbb{R}^{d}$, we have the following: $\forall i\in[n],\forall t\geq 0$, (1) $\mathbb{E}[\bm{g}_{i}(\bm{x},\bm{\xi}_{i}^{t})\mid\mathcal{F}_{t}]=\nabla f_{i}(\bm{x})$; (2) There exist $p\in(1,2]$, some constant $\sigma\geq 0$ such that $\mathbb{E}\big[\|\bm{g}_{i}(\bm{x},\bm{\xi}_{i}^{t})-\nabla f_{i}(\bm{x})\|^{p}\mid\mathcal{F}_{t}\big]\leq\sigma^{p}$; (3) The family $\{\bm{\xi}_{i}^{t}:\forall t\geq 0,i\in[n]\}$ of random samples is independent.
@@
Assumption 3 covers a broad class of heavy-tailed distributions, including Lévy’s $\alpha$-stable distributions, Student’s $t$-distributions, and Pareto distributions. Note that we do not assume noise symmetry as in [30], and when $p=2$, Assumption 3 reduces to the standard bounded variance condition commonly assumed in the literature.
@@
The nonnegative weight matrix $\bm{W}$, whose $(i,r)$-th component of $\bm{W}$, denoted as $w_{ir}$, is positive if and only if $(i,r)\in\mathcal{E}$ or $i=r$, is primitive and doubly stochastic, i.e., $\mathbf{1}_{n}\bm{W}=\mathbf{1}_{n}$ and $\bm{W}\mathbf{1}_{n}=\mathbf{1}_{n}$.
Assumption 4 is standard in the decentralized optimization literature [51], and it guarantees that there exists some nonnegative $\lambda$, i.e., spectral gap, such that
@@ 3 Algorithm Development: GT-NSGDm
We now describe the proposed Algorithm GT-NSGDm and discuss the intuition of its construction. We use $\bm{x}_{i}^{t}$ to denote the estimate of a stationary point for the global cost function $f$ at node $i$ and $t$-th iteration, and recall that $\bm{g}_{i}(\bm{x}_{i}^{t},\bm{\xi}_{i}^{t})$ denotes the corresponding stochastic gradient returned from a local first-order oracle. Motivated by the error-feedback approach in [30], which serves as a momentum-type of variance reduction after applying a nonlinear operator to handle heavy-tailed noise, we also employ local momentum variance reduction
@@
where $\beta\in[0,1)$ serves as the momentum coefficient. Then, we use gradient tracking [12] to handle heterogeneous local functions $\{f_{i}\}_{i=1}^{n}$. Specifically, we use an estimator $\bm{y}_{i}^{t}$ to track the global gradient
@@
It is known that gradient tracking helps eliminate the dependence on heterogeneity among local functions $\{f_{i}\}_{i=1}^{n}$, such as the requirement of bounded gradient similarity. Furthermore, similar to the approach in [27], which uses normalization to address heavy-tailed noise in centralized settings, we avoid applying normalization in the recursive updates of the local gradient estimator $\bm{v}_{i}^{t}$ in (1) and the global gradient tracker $\bm{y}_{i}^{t}$. Instead, normalization is applied only during the update of $\bm{x}_{i}^{t}$, with step size $\alpha$, and nonnegative mixing weights $\{w_{ir}\}$ where $w_{ir}>0$ only when $(i,r)\in\mathcal{E}$ or $i=r$,
@@
We combine the local updates (1)(2)(3) on node $i\in\mathcal{V}$ and call it GT-NSGDm, Gradient Tracking based Normalized Stochastic Gradient Descent with momentum. When taking $\beta=0$, this simplifies to momentum-free gradient tracking with normalization in step 3. However, our analysis shows that GT-NSGDm performs optimally for some $\beta\in(0,1)$, making GT-NSGDm a non-trivial and optimal algorithmic design for the considered problem class. We provide a tabular description for GT-NSGDm in Algorithm 1, where all $\{\bm{x}_{i}^{0}\}$ are initialized from the same point $\bar{\bm{x}}^{0}$ for simplicity.
@@
Then, the global average $\bar{\bm{x}}^{t}$ would update in the negative direction of the sum of normalized local gradients: $\bar{\bm{x}}^{t+1}=\bar{\bm{x}}^{t}-\frac{\alpha}{n}\sum_{r=1}^{n}\frac{\nabla f_{r}(\bm{x}_{r}^{t})}{\|\nabla f_{r}(\bm{x}_{r}^{t})\|}.$ Let for some $t,\forall r\in\mathcal{V},\bm{x}_{r}^{t}=\bm{x}_{*}=\mathop{\mathrm{arg\,min}}\sum_{i=1}^{n}f_{i}(\bm{x})$, i.e., all nodes hold the optimal global solution that $\sum_{r=1}^{n}\nabla f_{r}(\bm{x}_{*})=\mathbf{0}$. Since $\|\nabla f_{r}(\bm{x}_{*})\|$ can be different quantities for $r=1,\ldots,n$, due to function heterogeneity, then $\bar{\bm{x}}^{t+1}$ will move away from $\bm{x}_{*}$. Therefore, vanilla gradient normalization adds some intrinsic errors from heterogeneous local normalizations. By incorporating gradient tracking, we expect that $\bm{y}_{r}^{t}$ would converge to its global average $\bar{\bm{y}}^{t}$, and $\bar{\bm{y}}^{t}$ would converge to $(1/n)\sum_{i=r}^{n}\nabla f_{i}(\bm{x}_{r}^{t})$. In this way, $\bm{x}_{r}^{t}$ would move along the direction of the normalized sum of local gradients, and thus emulating the centralized setting.
@@
We next break the limitations of vanilla gradient normalization in Claim 1 by incorporating gradient tracking and momentum variance reduction. This enables the successful use of normalization to suppress heavy-tailed noise while maintaining optimal convergence despite the added nonlinearity.